# Isentropic processes for axisymmetric Black Holes

### v1

@@ Abstract
We demonstrate that the isentropic absorption of a classical charged test particle is classically forbidden for all 3+1 dimensional stationary, non-extremal, axisymmetric black holes in any diffeomorphism invariant theory of gravity. This result is derived purely from the near-horizon geometry and thermodynamic properties of the black hole spacetime, independent of the specific gravitational theory. We further consider the Kerr–Newman black hole in general relativity and analyze within the semiclassical WKB approximation, the conditions under which isentropic absorption may become allowed. Broader implications for the second law and extremality bounds are discussed.
@@ 1 Introduction
Beginning in the 1970s, several independent universal entropy bounds were proposed for arbitrary physical systems. Among the earliest was the Bekenstein bound, formulated using semiclassical arguments to suggest that the maximum entropy of an object with finite energy and spatial extent is bounded from above [1, 2, 3, 4]. One can also interpret it as an upper bound on the information contained by a system with a finite energy and spatial extent [2]. In this context, [5] examined the constraints on a channel’s communication capacity imposed by the Bekenstein bound. Further, [6] explored the validity of the generalized Bekenstein bound for the bound states of interacting quantum fields and found that, if interactions are properly included, the entropy bound holds valid. [7] used the unitarity of scattering amplitudes to show that for a self-sustained quantum field object of size $R$, the universal entropy bound can be expressed as $S\leq 1/\alpha$, where the running coupling $\alpha$ is evaluated at the scale $1/R$. Black holes are known to saturate the entropy bound with their entropy expressible as $S=1/\alpha$ where $\alpha=G/R^{d-2}$, and $G$ is Newton’s gravitational constant in dimensions $d$ [7, 8]. Various other derivations and formulations of the bound have been proposed in the literature. For instance, [9] showed from a holographic perspective that the Bekenstein bound holds for any strength of the gravitational coupling under certain assumptions. There are also several alternate or generalized forms of the entropy bounds, such as the causal entropy bound [10], the Unruh-Wald bound [11, 12], generalized Bekenstein bound [13], covariant entropy bound [14, 15], the generalized covariant entropy bound [16, 17], etc.
There are also several interpretations of the Bekenstein bound. For instance [18] illustrated that the Bekenstein bound naturally arises as a consequence of the Pauli exclusion principle when one considers the fundamental degrees of freedom of the black hole to be finite in number and fermionic in nature. Detailed study of the Bekenstein bound for the black holes’ fermionic degrees of freedom can be found in [19, 20]. The Bekenstein bound imposes significant constraints on the information content and thermodynamic behavior of a black hole over its lifetime. [21] applied the Bekenstein entropy bound to analyze the re-distribution of the fundamental degrees of freedom of generic physical systems and found an average information loss due to the entanglement between geometry and fields. Further, [22] used the entropy bound for constraining the information capacity of black hole remnants and concluded that the black hole remnants, as a final stage of black hole evaporation, might not be a viable resolution of the information loss paradox. The energy density of a quantum field can be both positive and negative, provided that the negative energy density in some region of space is balanced by a positive energy density in other regions. [23] illustrated that the Bekenstein bound requires the region containing negative energy density not to be arbitrarily far from the region with positive energy density; that is, there must be a localization of negative energy. One can refer [24, 25, 26] for the consequences of the negative energy density in the time evolution of a black hole.
The black hole information loss paradox suggests that certain fundamental principles of modern physics, namely, unitarity, the applicability of local quantum field theory in describing black hole evaporation, general covariance, the Bekenstein-Hawking entropy relation with the Boltzman entropy of the black hole, and the existence of a semi-classical observer capable of retrieving information from Hawking radiation cannot all simultaneously hold [27, 28, 29, 30]. Further, to resolve such issues, one perhaps needs to take non-perturbative effects into account. These non-perturbative effects may be less dominant than the semiclassical contributions during the initial stages of evaporation; however they can accumulate and provide a significant contribution over the lifetime of the black hole. The relevance of these effects depends upon several factors such as the dimensions of spacetime [31], the memory burden effect [32, 33], fluctuations in the gravitational degrees of freedom [34], existence of critical regimes and instabilities [33, 35], etc. The identification that the Bekenstein-Hawking entropy of a black hole is proportional to its Boltzmann entropy [2] relies on the semiclassical approximation. However, it has been argued in [28] that the non-perturbative effects can lead to the breakdown of the semiclassical approximation and can allow the Boltzmann entropy inside a black hole to be greater than its Bekenstein-Hawking entropy. In particular, [28] analyzed the isentropic absorption of charged particles by a Reisner-Nordstrom black hole within WKB approximation and proposed that such a process could be a non-perturbative way through which entropy bound can be exceeded. In this article, we consider the isentropic absorption of charged particles by a stationary axi-symmetric black hole within WKB approximation processes and discuss implications in the context of the entropy bound.
In section 2, we discuss a general stationary nonextremal axisymmetric black hole spacetime and show that classically it is not possible to absorb a charged particle without changing the black hole’s entropy. There exists a classically forbidden region outside the black hole for such an absorption. However, quantum mechanically, it is possible for certain values of the test particle and the black hole parameters that the test particle tunnel through this potential barrier and get absorbed by the black hole. In section 3 we use the WKB approximation to compute the tunneling probability. Finally in section 4 we discuss the implications of this type of absorption. We use units where $\hbar=c=k_{B}=G=1$, and adopt a mostly positive metric signature ($+$)(-$+$)( - + + + ).
@@ 2 Charged particle in stationary axisymmetric spacetime
The metric functions are assumed to be smooth, continuous, atleast $C^{2}$ differentiable, independent of the $t$ and $\phi$ coordinates and thus adapted to the Killing symmetries of stationarity, $\partial_{t}$ and axi-symmetry $\partial_{\phi}$. The spacetime is assumed to be circular and asymptotically flat. Further, the black hole spacetime is assumed to possess a Killing horizon at $r=r_{+}$ such that $g^{rr}(r_{+},\theta)=0$. The black hole is described by its mass $M$, charge $Q$ and angular momemtum $J$.
@@
$-\epsilon=g_{tt}(r,\theta)\bigg{(}\frac{dt}{d\lambda}\bigg{)}^{2}+2g_{t\phi}(r,\theta)\frac{dt}{d\lambda}\frac{d\phi}{d\lambda}+g_{\phi\phi}(r,\theta)\bigg{(}\frac{d\phi}{d\lambda}\bigg{)}^{2}+g_{rr}(r,\theta)\bigg{(}\frac{dr}{d\lambda}\bigg{)}^{2}+g_{\theta\theta}(r,\theta)\bigg{(}\frac{d\theta}{d\lambda}\bigg{)}^{2},$ (7)
@@
The first and second terms in equation (8) are positive which implies that the third term denoted by $V_{eff}$ must be negative for the expression to be valid for a classical test particle. Further, we stress that Eqs.(8)-(9) are valid for any diffeomorphism invariant theory of gravity that admits a black hole solution of the form described in (1).
@@ 2.1 Isentropic absorption of the charged particle
The change in mass $dM$, charge $dQ$, and angular momentum $dJ$ of the blackhole, when the charged particle enters the horizon, is given by $E$, $q$, and $L$ respectively. Further, the change of electromagnetic potential energy is given by $q\Phi_{e}=qA_{t}-\Omega qA_{\phi}$. Therefore, the above equation gives
$E-qA_{t}+q\Omega A_{\phi}-\Omega L=0,$ (11)
which is $E^{\prime}=\Omega L^{\prime}$, at the event horizon. Therefore, the effective potential at the horizon can be written as
$V_{eff}(r=r_{+},\theta)=-\frac{g^{rr}L^{\prime 2}(\Omega^{2}g_{\phi\phi}+g_{tt}+2g_{t\phi}\Omega)}{m^{2}D}\bigg{|}_{r=r+}+\epsilon g^{rr}\bigg{|}_{r=r+}$ (12)
The spacetime under consideration is stationary, so the event horizon will be a Killing horizon [39, 40]. One can consider the Killing horizon to be generated by the Killing vector
@@
The effective potential at the horizon in terms of $\chi$ can be rewritten as
@@
where prime on $D$ and the metric components denote the derivative with respect to $r$. Evaluating the above expression on the horizon for an isentropic process with $E^{\prime}=\Omega L^{\prime}$ one gets,
@@
since $g^{\prime rr}>0$ and $g_{\phi\phi}>0$ at the horizon. Since $V_{eff}(r_{+},\theta)=0$ and $dV_{eff}(r_{+},\theta)/dr>0$, we necessarily have $V_{eff}>0$ outside the horizon and hence, classically it is not possible to absorb the charged particle isentropically.
@@ 2.2 Kerr-Newman black holes in general relativity
Here, $a$ is the specific angular momentum, $M>0$ is the ADM mass, and $Q$ is the electric charge of the blackhole. Assuming a vanishing magnetic charge, the electromagnetic 4-vector potential of the black hole is given by $A_{\mu}=(-Qr/\Sigma,0,0,Qar\sin^{2}\theta/\Sigma)$. So, the sign convention is such that a positive value of $Q$ means a positively charged black hole. One can see by substituting the metric Eq.(2.2) in equations (12) and (16) that for the isentropic absorption of the charged particle
@@
by direct calculation. Therefore, the isentropic absorption of the charged test particle is not possible classically.
@@ 3 Quantum tunneling
We saw in the previous section that the effective potential vanishes at the event horizon and that its derivative becomes positive for the isentropic absorption. We plot the effective potential in Fig.[1], from which it can be seen that for certain values of charge, mass, and angular momentum, the effective potential can become negative again at some distance from the event horizon. The region between the event horizon at $r=r_{+}$ and the point outside, say $r=r_{2}>r_{+}$, where the effective potential vanishes again, is classically forbidden. We see from Fig. [1] that increasing the angular momentum of the particle increases both the width of the classically forbidden region and the height of $V_{eff}$. However, the increase in charge reduces both. Further, we see that the change of mass of the test particle doesn’t change the width of the classically forbidden region, but the height of the potential barrier experienced changes. So, the massive charged test particles with low angular momentum experience a small, but nonzero, potential for the isentropic absorption.
The tunneling probability through the classically forbidden region between $r_{+}$ and $r_{2}$ in the WKB approximation is given by [28]
$\Gamma\approx e^{-2S_{E}},$ (25)
where $S_{E}$ is the Euclidean action given by
[…]
@@
Here, we have assumed $d\theta/d\lambda=0$ for simplicity and used Eq.(8) with $\lambda_{E}=i\lambda$. We choose negative root in $dr/d\lambda_{E}=\pm\sqrt{V_{eff}}$ because the particle is being absorved by the black hole. It’s difficult to evaluate the integral in Eq. (29) analytically, so we compute it numerically and plot the corresponding Euclidean action for a real value of turning points $r_{+}$ and $r_{2}$. The tunneling probability in Eq.(25) decreases (increases) as the Euclidean action increases (decreases), so Fig.[2]-[3] effectively describe the tunneling probability in the WKB approximation.
@@
One can notice from Fig.[2] that the Euclidean action increases with an increase in the angular momentum of the particle for all combinations of parameters. Therefore, the tunneling probability for low angular momentum particles is large in comparison with the higher angular momentum particles. One can understand it as the effective potential as well as the width of the classically forbidden region also increases with an increase in the angular momentum of the test particle. Furthermore, the increase in the magnitude of charge $Q$ of the blackhole also reduces the Euclidean action. However, the increase in mass of the blackhole increases the Euclidean action. The increase of the specific angular momentum $a$ of the black hole decreases the effective action for the positive $L$ particles, but increases the effective action for the negative $L$ particles. Therefore, a low mass, high specific angular momentum, and large charge black hole has a higher tunneling probability for a test charged particle. The Euclidean action is smaller for a larger polar angle. Therefore, the isentropic absorption is more probable near the equator of the black hole.
@@
In Fig.[3], we see that the increase of the size of the black hole, determined by its outer horizon $r_{+}$, increases the Euclidean action for all combinations of parameters. Therefore, smaller black holes possess a greater tunneling probability than larger black holes. This suggests that the smaller black holes are relatively less classical as they can absorb charged particles isentropically and are prone to violate the entropy bound more efficiently than a large black hole. Furthermore, we notice that increasing the mass and the magnitude of angular momentum of the test particle increases the Euclidean action for a fixed size $r_{+}$ of the black hole, whereas an increase in the magnitude of the charge of the test particles decreases the Euclidean action. Therefore, lighter particles with small angular momentum, but large charge, have a large tunneling probability. Further, the tunneling probability is larger near the equator for a fixed $r_{+}$.

### v2

@@ Abstract
We demonstrate that the isentropic absorption of a classical charged test particle is classically forbidden for all (3+1)-dimensional stationary, non-extremal, axisymmetric black holes in any diffeomorphism invariant theory of gravity. This result is derived purely from the near-horizon geometry and thermodynamic properties of the black hole spacetime, independent of the specific gravitational theory. We further consider the Kerr–Newman black hole in general relativity and analyse, using the quantum tunnelling approach, the conditions under which isentropic absorption may be allowed. Broader implications for the second law and extremality bounds are discussed.
@@ 1 Introduction
Beginning in the 1970s, several independent universal entropy bounds were proposed for arbitrary physical systems. Among the earliest was the Bekenstein bound, formulated using semiclassical arguments to suggest that the maximum entropy of an object with finite energy and spatial extent is bounded from above [1, 2, 3, 4]. This result can also be interpreted as an upper limit on the information content of such a system [2]. Since then, the Bekenstein bound has been explored and extended in a variety of directions. In particular, [5] examined how it constrains the communication capacity of quantum channels, while [6] analysed its validity for bound states of interacting fields and showed that, once interactions are properly included, the bound remains intact. From a scattering perspective, [7] demonstrated that for a self-sustained quantum field object of size $R$, unitarity implies $S\leq 1/\alpha$, where $\alpha$ is the running coupling evaluated at the scale $1/R$. Black holes, in turn, are known to saturate this bound, with their entropy taking the form $S=1/\alpha$, where $\alpha=G/R^{d-2}$ and $G$ is Newton’s constant in $d$ dimensions [7, 8]. Additional derivations further reinforce the universality of this limit. For instance, [9] showed from a holographic perspective that the Bekenstein bound holds for any strength of the gravitational coupling under certain assumptions, while a wide variety of generalised forms have also been proposed, including the causal entropy bound [10], the Unruh–Wald bound [11, 12], generalized Bekenstein bounds [13], covariant entropy bounds [14, 15], and their generalized covariant versions [16, 17].
Alongside various formulations mentioned in the previous paragraph, different interpretations of the Bekenstein bound have also been explored. We recall some of them below. For instance, [18] argued that it arises naturally from the Pauli exclusion principle if the fundamental degrees of freedom of a black hole are finite and fermionic. Related studies of such fermionic degrees of freedom are found in [19, 20]. Other works emphasise the implications of the bound for information dynamics: [21] showed that the entropy bound governs the redistribution of degrees of freedom in generic systems and implies an average information loss due to entanglement between geometry and fields, while [22] used it to constrain the information capacity of black hole remnants, arguing against remnants as a resolution of the information paradox. The interplay between entropy bounds and negative energy has also been highlighted. Since quantum fields admit regions of negative energy density provided they are compensated elsewhere, [23] showed that the Bekenstein bound requires such negative-energy regions to be localised near positive-energy ones. The consequences of negative energy densities for black hole dynamics have been further examined in [24, 25, 26]
All the considerations highlighted above are directly relevant for the black hole information loss paradox, which reveal tensions among fundamental principles: unitarity, local quantum field theory, general covariance, the identification of Bekenstein–Hawking entropy with Boltzmann entropy, and the existence of a semiclassical observer retrieving information from Hawking radiation cannot all simultaneously hold [27, 28, 29, 30]. One possible resolution is that non-perturbative effects—though negligible in the early stages of evaporation—can accumulate and become significant over the black hole’s lifetime. Such effects may depend on spacetime dimensionality [31], memory burden phenomena [32, 33], fluctuations in gravitational degrees of freedom [34], or critical instabilities [33, 35]. Although the non-perturbative mechanisms operate through distinct physical channels, they share a common capacity to accumulate effects gradually and ultimately perhaps dominate the entropy dynamics, thereby breaking the semiclassical predictions. Importantly, while the identification of Bekenstein–Hawking entropy with Boltzmann entropy is semiclassical [2], it has been argued in [28] that non-perturbative contributions can break this equivalence and allow the Boltzmann entropy to exceed the Bekenstein–Hawking value. In particular, [28] analysed the isentropic absorption of charged particles by a Reissner–Nordström black hole within the WKB approximation, suggesting that such a process may serve as a non-perturbative channel for exceeding entropy bounds.
Motivated by the findings mentioned above, in the present article, we extend the analysis of the isentropic absorption of charged particles to stationary, axisymmetric black holes. We first demonstrate in Section 2 that, for non-extremal axisymmetric black holes, the absorption of a charged particle without changing the Bekenstein–Hawking entropy is classically forbidden: a classically forbidden region exists outside the horizon. The result of the existence of the classically forbidden region is rooted in the near-horizon geometry and thermodynamic laws–it holds across all diffeomorphism-invariant gravitational theories. We then turn, in Section 3, to the quantum regime, where tunnelling across the barrier becomes possible. Using the saddle-point approximation, we compute tunnelling probabilities and show how they depend on both the particle’s and the black hole’s parameters. Finally, Section 4, delves into the broader significance of these processes: small, highly charged black holes emerge as more susceptible to isentropic absorption, potentially allowing entanglement (hence Boltzmann) entropy to grow—raising subtle questions regarding entropy bounds and information accumulation. We use units $\hbar=c=k_{B}=G=1$ and adopt a mostly positive metric signature (+++)(-+++).
@@ 2 Charged particle in stationary axisymmetric spacetime
The metric functions are assumed to be smooth and continuous, at least $C^{2}$-differentiable, independent of the $t$ and $\phi$ coordinates, and thus adapted to the Killing symmetries of stationarity ($\partial_{t}$) and axisymmetry ($\partial_{\phi}$). The spacetime is assumed to be circular and asymptotically flat. Further, the black hole spacetime is assumed to possess a Killing horizon at $r=r_{+}$ such that $g^{rr}(r_{+},\theta)=0$. The black hole is described by its mass $M$, charge $Q$, and angular momentum $J$.
@@
$\epsilon=g_{tt}(r,\theta)\bigg(\frac{dt}{d\lambda}\bigg)^{2}+2g_{t\phi}(r,\theta)\frac{dt}{d\lambda}\frac{d\phi}{d\lambda}+g_{\phi\phi}(r,\theta)\bigg(\frac{d\phi}{d\lambda}\bigg)^{2}+g_{rr}(r,\theta)\bigg(\frac{dr}{d\lambda}\bigg)^{2}+g_{\theta\theta}(r,\theta)\bigg(\frac{d\theta}{d\lambda}\bigg)^{2},$ (7)
@@
The first and second terms in equation (8) are positive, which implies that the third term, denoted by $V_{eff}$, must be negative for the expression to be valid for a classical test particle. Further, we stress that Eqs.(8)-(9) are valid for any diffeomorphism invariant theory of gravity that admits a black hole solution of the form described in (1).
@@ 2.1 Isentropic absorption of the charged particle
The change in mass $dM$, charge $dQ$, and angular momentum $dJ$ of the black hole, when the charged particle enters the horizon, is given by $E$, $q$, and $L$, respectively. The spacetime under consideration is stationary, so the event horizon will be a Killing horizon [39, 40]. One can consider the Killing horizon to be generated by the Killing vector
@@
Here, $\Omega_{H}$ denotes the angular velocity evaluated at the outer horizon of the black hole, which is interpreted as the angular velocity of the black hole itself. The electric potential on the horizon is defined by $\Phi_{e}\equiv-\,\chi^{\mu}A_{\mu}\big|_{r_{+}}$ [41]. Using our convention in the Lorentz gauge for the gauge field, $A_{\mu}=(-A_{t},0,0,A_{\phi})$, the contraction gives $\Phi_{e}=A_{t}-\Omega_{H}A_{\phi},$ so that the change in electromagnetic potential energy of the black hole due to absorption of a particle of charge $q$ is $q\Phi_{e}=qA_{t}-\Omega_{H}qA_{\phi}$. Therefore, the above equation (10) gives
$E-qA_{t}+q\Omega_{H}A_{\phi}-\Omega_{H}L=0,$ (12)
which is $E^{\prime}=\Omega_{H}L^{\prime}$, at the event horizon. Therefore, the effective potential at the horizon can be written as
$V_{eff}(r=r_{+},\theta)=-\frac{g^{rr}L^{\prime 2}(\Omega_{H}^{2}g_{\phi\phi}+g_{tt}+2g_{t\phi}\Omega_{H})}{m^{2}D}\bigg|_{r=r+}+\epsilon g^{rr}\bigg|_{r=r+}$ (13)
The above expression for the effective potential at the horizon in terms of the Killing vector $\chi^{\mu}$ can be rewritten as
@@
where prime on $D$ and the metric components denote the derivative with respect to $r$. Evaluating the above expression on the horizon for an isentropic process with $E^{\prime}=\Omega L^{\prime}$, one gets,
@@
since $g^{\prime rr}>0$ and $g_{\phi\phi}>0$ at the horizon. Since $V_{eff}(r_{+},\theta)=0$ and $dV_{eff}(r_{+},\theta)/dr>0$, we necessarily have $V_{eff}>0$ outside the horizon and hence, classically, it is not possible to absorb the charged particle isentropically.
@@ 2.2 Kerr-Newman black holes in general relativity
Here, $a$ is the specific angular momentum, $M>0$ is the ADM mass, and $Q$ is the electric charge of the black hole. Assuming a vanishing magnetic charge, the electromagnetic 4-vector potential of the black hole is given by $A_{\mu}=(-Qr/\Sigma,0,0,Qar\sin^{2}\theta/\Sigma)$. So, the sign convention is such that a positive value of $Q$ means a positively charged black hole. By substituting the metric in Eq.(2.2) into Eqs.(13) and (16), one obtains, respectively, for the isentropic absorption of a charged particle,
@@
The positivity of $dV_{eff}/dr$ at $r=r_{+}$ in Eq. (24) follows from the fact that the term inside the square brackets is strictly positive, while the prefactor is also positive, since $Mr_{+}=M^{2}+M\sqrt{M^{2}-a^{2}-Q^{2}},$ and the non-extremality condition requires $M^{2}>a^{2}+Q^{2}$. Therefore, the isentropic absorption of the charged test particle is not possible classically.
@@ 3 Quantum tunneling
We saw in the previous section that the effective potential vanishes at the event horizon and that its derivative becomes positive for the isentropic absorption. We plot the effective potential in Fig.[1], from which it can be seen that for certain values of charge, mass, and angular momentum, the effective potential can become negative again at some distance from the event horizon. The region between the event horizon at $r=r_{+}$ and the point outside, say $r=r_{2}>r_{+}$, where the effective potential vanishes again, is classically forbidden. We see from Fig. [1] that increasing the angular momentum of the particle increases both the width of the classically forbidden region and the height of $V_{eff}$. However, the increase in charge reduces both. Further, we see that the change of mass of the test particle does not change the width of the classically forbidden region, but the height of the potential barrier experienced changes. So, the massive charged test particles with low angular momentum experience a small, but nonzero, potential for the isentropic absorption.
Feynman’s path integral formulation expresses the transition amplitude for a radial degree of freedom $r(\lambda)$, evolving in an effective potential $V_{eff}$, as
$K\;\propto\;\int\mathcal{D}r\;\exp\!\left(\tfrac{i}{\hbar}S\right),$ (25)
where $S$ is the action. The quantum dynamics is found by summing over all admissible trajectories $r(\lambda)$, each weighted by the phase factor $\exp(iS/\hbar)$ determined by the action. In the classically forbidden region, where $V_{eff}>0$, the path integral is dominated by the steepest-descent contour obtained via a Wick rotation to imaginary time, $\lambda\to-i\lambda_{E}$. Under this transformation, the action becomes Euclidean, $S\to iS_{E}$, and the amplitude reduces to
$K\;\sim\;e^{-S_{E}/\hbar},\qquad S_{E}=\int d\lambda_{E}\,L_{E}\,$, (26)
where $L_{E}$ is the Euclideanized Lagrangian. The dominant contribution comes from the Euclidean classical trajectory—the solution of the Euler–Lagrange equations in imaginary time—connecting the turning points of the potential barrier. Since the amplitude acquires a factor $e^{-S_{E}/\hbar}$, the tunnelling probability is obtained by taking the modulus squared of this amplitude,
$\Gamma\;\propto\;|K|^{2}\;\sim\;e^{-2S_{E}/\hbar}.$ (27)
So, setting $\hbar=1$, the probability through the forbidden region $r_{2},r_{+}]$ can be expressed as [28]
$\Gamma\;\approx\;e^{-2S_{E}}.$ (28)
Here, the Euclidean action $S_{E}$ is given by
$S_{E}$ $=\int_{r_{2}}^{r_{+}}L_{E}d\lambda_{E}$ (29)
$=\int_{r_{2}}^{r_{+}}\frac{1}{dr/d\lambda_{E}}dr$ (30)
$=-\int_{r_{2}}^{r_{+}}\frac{1}{\sqrt{V_{eff}(r)}}dr$ (31)
$=\int_{r_{+}}^{r_{2}}\frac{1}{\sqrt{V_{eff}(r)}}dr.$ (32)
@@
Here, we have assumed $d\theta/d\lambda=0$ for simplicity and have used Eq. (8). We choose the negative root in $dr/d\lambda_{E}=\pm\sqrt{V_{eff}}$ because the particle is being absorbed by the black hole. It is difficult to evaluate the integral in Eq. (32) analytically, so we compute it numerically and plot the corresponding Euclidean action for a real value of turning points $r_{+}$ and $r_{2}$. The tunnelling probability in Eq.(28) decreases (increases) as the Euclidean action increases (decreases), so Fig.[2]-[3] effectively describe the tunnelling probability in the approximation discussed above.
@@
One can notice from Fig.[2] that the Euclidean action increases with an increase in the angular momentum of the particle for all combinations of parameters. Therefore, the tunnelling probability for low angular momentum particles is large in comparison with the higher angular momentum particles. One can understand it as the effective potential, as well as the width of the classically forbidden region, increasing with an increase in the angular momentum of the test particle. Furthermore, the increase in the magnitude of charge $Q$ of the black hole reduces the Euclidean action. However, the increase in mass of the black hole increases the Euclidean action. The increase of the specific angular momentum $a$ of the black hole decreases the effective action for the positive $L$ particles, but increases the effective action for the negative $L$ particles. Therefore, a low mass, high specific angular momentum, and large charge black hole has a higher tunnelling probability for a test charged particle. The Euclidean action is smaller for a larger polar angle. Therefore, the isentropic absorption is more probable near the equator of the black hole.
@@
In Fig.[3], we see that the increase in the size of the black hole, determined by its outer horizon $r_{+}$, increases the Euclidean action for all combinations of parameters. Therefore, smaller black holes have a greater tunnelling probability than larger black holes. This suggests that the smaller black holes are relatively less classical as they can absorb charged particles isentropically and are prone to violate the entropy bound more efficiently than a large black hole. Furthermore, we notice that increasing the mass and the magnitude of angular momentum of the test particle increases the Euclidean action for a fixed size $r_{+}$ of the black hole, whereas an increase in the magnitude of the charge of the test particles decreases the Euclidean action. Therefore, lighter particles with small angular momentum, but large charge, have a large tunnelling probability. Further, the tunnelling probability is larger near the equator for a fixed $r_{+}$.